# Dark Matter Search Results from 4.2 Tonne-Years of Exposure of the LUX-ZEPLIN (LZ) Experiment J. Aalbers

### v2

@@ Appendix A Appendix - Spin Dependent Results
de Laeter et al. [2003a]J. R. de Laeter, J. K. Böhlke, P. D. Bièvre, H. Hidaka, H. S. Peiser, K. J. R. Rosman, and P. D. P. Taylor, Pure and Applied Chemistry 75, 683 (2003a).
@@
Szydagis et al. [2022]M. Szydagis, J. Balajthy, G. Block, J. Brodsky, E. Brown, J. Cutter, S. Farrell, J. Huang, E. Kozlova, C. Liebenthal, et al., arXiv:2211.10726 (2022).
de Laeter et al. [2003b]J. R. de Laeter, J. K. Böhlke, P. D. Bièvre, H. Hidaka, H. S. Peiser, K. J. R. Rosman, and P. D. P. Taylor, Pure Appl. Chem. 75, 683 (2003b).
@@ Appendix D WS2024-Only Limit
Figure S2 shows the SI WIMP-nucleon cross section upper limit obtained using the data and models pertaining to the WS2024, 3.3 tonne year exposure only.

### v3

@@ Appendix A Appendix - Spin Dependent Results
Szydagis et al. [2025]M. Szydagis, J. Balajthy, G. Block, J. Brodsky, E. Brown, J. Cutter, S. Farrell, J. Huang, E. Kozlova, C. Liebenthal, et al., Front. Detect. Sci. Technol. 2 (2025).
@@
de Laeter et al. [2003]J. R. de Laeter, J. K. Böhlke, P. D. Bièvre, H. Hidaka, H. S. Peiser, K. J. R. Rosman, and P. D. P. Taylor, Pure and Applied Chemistry 75, 683 (2003).
@@ Appendix D WS2024-Only Limit
Figure S2 shows the SI WIMP-nucleon cross section upper limit obtained using the data and models pertaining to the WS2024, 3.3 tonne-year exposure only.
@@ Appendix E Limits in Terms of Number of WIMP Events
Appendix F Parameters of the Salting Model
As discussed in the main text, the salt is drawn from a distribution consisting of the sum of a WIMP-like exponential and a flat spectrum, as shown in Fig. S4. The parameter sets the maximum energy of the flat distribution and is drawn from a Gaussian distribution with mean keV and sigma keV. The parameter is chosen to mimic the recoil spectrum induced by a WIMP with mass , which is in turn drawn from a Gaussian with mean = 80 GeV/c2 and width = 20 GeV/c2. The normalization is drawn from a uniform distribution that spans from zero up to a rate that would be consistent with the upper limits from Ref. [8].
Figure S4: Distribution from which the salt events are drawn. The parameters are described in the text.
The actual normalization, exponential, and flat spectrum parameters used are still hidden to the collaboration as unveiling them would harm bias mitigation of ongoing analyses. The allowed ranges were chosen to ensure that salt events can plausibly span the WIMP-search energy range, and also reach the highest recoil energies that might be used in a future higher-energy analysis, such as for effective field theories of dark matter.